3x-11y=20
2x=4y+10

Answer by JulietG(1812)

3x-11y=20
2x=4y+10
.
Take a look at the second equation. What do you notice?
We can see that x can easily be found.
2x=4y+10
Divide each side by 2
X = 2Y + 5
Now, take the value of X and put it into the first equation.
3x-11y=20
3(2Y+5)-11y=20
Distribute
6Y+5-11y=20
Subtract 5 from each side
6Y-11Y=15
-5Y = 15
Divide each side by -5
Y = -3
.
x = 2Y + 5
x = -6 + 5
x = -1
.
Let's prove it by putting it into the equation above.
2x=4y+10
2(-1)=4(-3)+10
-2 = 12+10
-2 = -2
Success!